Operations Report – November 29th

Crew 183 + Operations Report 29NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 5:30 pm

Solar— SOC  91% (Before generator is run at night)

Diesel –  50 %

Propane –  100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons.

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 3.5

Notes on ATVs: nothing to report

Deimos rover used:  no

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments:  Propane was delivered today.  A leak was
found in the line of the small tank at the director’s house, and it
was fixed.  We began work on the second rover today but then decided
to go to storage and see it we could find better tools.  We found a
socket set that was the right size!  We will get back on it tomorrow.
We also brought a 55 gallon drum from town.  Tomorrow I will pound
holes in it and set it up as a burn barrel.  I will also need to clean
up the old burn barrel area as non-burnable garbage was burned and it
will need to be picked up and taken for disposal.  Lucky me!

Summary of internet:  not known

Summary of suits and radios: nothing to report

Summary of Hab operations: Three new smoke detectors arrived today.

Summary of GreenHab operations: Door was opened this afternoon.  Two
gallons of water was used in the new planters.  I did not water the
plants today.

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Thanks for all of
the help with the rover installation question last night.  No
questions tonight.